OpenStudy (anonymous):

Find the x- and y-intercepts for the equation. -x – 3y = 9

OpenStudy (anonymous):

See for finding x intercept just put y = 0 and for finding y intercept just put x = 0 can you do that??

OpenStudy (anonymous):

@cheriekeeney getting or not???

OpenStudy (anonymous):

sorry i stepped away for a min

OpenStudy (anonymous):

and yes I can get that what i don't understand is where 9 comes from, everything i've tried comes up with -9

OpenStudy (anonymous):

Put x = 0 -3y = 9 y = -3

OpenStudy (anonymous):

Put y = 0 -x = 9 x = -9

OpenStudy (anonymous):

ok that is for x what about y that is where i was getting stuck

OpenStudy (anonymous):

I have found y above..

OpenStudy (anonymous):

so why is 0 for both sides???

OpenStudy (anonymous):

Where???

OpenStudy (anonymous):

See, firstly put y = 0 and find x...

OpenStudy (anonymous):

i'm sorry i was doing it the way someone showed me find out x first isolate y then find y and isolate x

OpenStudy (anonymous):

No... You can put directly..

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

Getting???

OpenStudy (anonymous):

yes thank you

OpenStudy (anonymous):

Welcome dear...